OpenStudy (anonymous):

Adanna had a glass cylinder with a diameter of 23 cm and a height of 10 cm. She placed 240 marbles, each with a diameter of 1.2 cm, into the cylinder. She filled the rest of the cylinder with colored water.How many cubic centimeters of colored water did she use? Use 3.14 to approximate pi and round your answer to one decimal place.

OpenStudy (welshfella):

first calculate the volume of the cylinder V = 3.14*r^2*h where h = radius and h = height

OpenStudy (welshfella):

the the volume of the 240 marbles volume of 1 marble = (4/3)*3.14*r^3 where r = radius of the marble

OpenStudy (welshfella):

you are given the diameters radius = half of the diameter

OpenStudy (welshfella):

* first post should be r = radius

OpenStudy (anonymous):

I followed all the steps and got 3 different answers "1.3939.6, 3935.7, and 3935.6."

OpenStudy (welshfella):

what did you get for the volume of the glass cylinder?

OpenStudy (welshfella):

volume of the water = volume of cylinder - volume of the 240 marbles

OpenStudy (anonymous):

for the volume I got4152.65

OpenStudy (welshfella):

thats right

OpenStudy (welshfella):

and the volume of the 240 marbles ?

OpenStudy (anonymous):

I got 217.0368

OpenStudy (welshfella):

yes thats right too

OpenStudy (welshfella):

I got 3935.6 for vol of water